A new car is purchased for 19900 dollars. The value of the car depreciates at 7.5% per year. What will the value of the car be, to the nearest cent, after 8 years?


1
Expert's answer
2021-06-30T10:20:07-0400

Let the price at which the car was purchased be P and the depreciated price in 8 years be Pd

Depreciated price is given by the formula below

"P_d=P(1-r)^n"

r=7.5%=0.075

P=$19900

n=8

"P_d=19900(1-0.075)^8\\
P_d=19900(0.925)^8\\
P_d=10666.64\\"

Value of the car after 8 years is $10666.64